OpenStudy (anonymous):

A doctor is doing a study of the height of children who eat their vegetables. He recorded the height of the children each year. The first year, the child was 36 inches. The second year, the child was 40 inches. The third year, the child was 44 inches. Write a function rule to represent the height of the child as an arithmetic sequence. How tall will the child be after 6 years?

OpenStudy (helder_edwin):

\[\large a_1=36 \text{in}\] \[\large a_2=40 \text{in} \] \[\large a_3=44\text{in} \] do u see the relation between each one and the next one?

OpenStudy (anonymous):

yess i do so is that the answer or are you explaining it

OpenStudy (helder_edwin):

each element (but the first one) is the previous one plus 4. do u see that?

OpenStudy (anonymous):

yes i see im starting to undertand it

OpenStudy (helder_edwin):

ok so \[\large a_i=a_{i-1}+4 \] for \(i>1\).

OpenStudy (anonymous):

okay so now i get it !! so easy thank you for your help

OpenStudy (helder_edwin):

u r welcome

OpenStudy (anonymous):

can you help with some more?

OpenStudy (helder_edwin):

sure

OpenStudy (anonymous):

Lee earns $6.25 per hour mowing lawns. a. Write a rule to describe how the amount of money m earned is a function of the number of hours h spent mowing lawns. b. How much does Lee earn if she works 5 hours and 15 minutes?

OpenStudy (helder_edwin):

1 hour 6.25 USD 2 hours \(6.25+6.25=\color{red}{2}\times6.25\) USD 3 hours \(6.25+.25+6.25=\color{red}{3}\times6.25\) USD so n hours ...

OpenStudy (anonymous):

so the total of n hours is ?

OpenStudy (helder_edwin):

can u figure out?

OpenStudy (anonymous):

39. 375 ?

OpenStudy (anonymous):

18.75 ?

OpenStudy (helder_edwin):

if u work 3 hours u get paid 3 times 6.25. so if u work h hours u get paid \[\large m=6.25\times h \]

OpenStudy (anonymous):

i understand it its goood perfect

OpenStudy (anonymous):

i wjust worked it out on paper

OpenStudy (helder_edwin):

15 min = 1/4 hour so for 5 hours and 15 min u get paid \[\large m=6.25\times(5+1/4)=\$32.8125 \]

OpenStudy (anonymous):

thats the exact samething i got thanks or helping me understand( :

OpenStudy (helder_edwin):

u r welcome